# Live Video Captioning

### v1

@@ Abstract
Dense video captioning is the task that involves the detection and description of events within video sequences. While traditional approaches focus on offline solutions where the entire video of analysis is available for the captioning model, in this work we introduce a paradigm shift towards Live Video Captioning (LVC). In LVC, dense video captioning models must generate captions for video streams in an online manner, facing important constraints such as having to work with partial observations of the video, the need for temporal anticipation and, of course, ensuring ideally a real-time response. In this work we formally introduce the novel problem of LVC and propose new evaluation metrics tailored for the online scenario, demonstrating their superiority over traditional metrics. We also propose an LVC model integrating deformable transformers and temporal filtering to address the LVC new challenges. Experimental evaluations on the ActivityNet Captions dataset validate the effectiveness of our approach, highlighting its performance in LVC compared to state-of-the-art offline methods. Results of our model as well as an evaluation kit with the novel metrics integrated are made publicly available to encourage further research on LVC.
Index Terms:
@@
dense video captioning, live video captioning, online, transformers, artificial intelligence, computer vision.
I Introduction
@@
AS a growing field within video understanding, video captioning has gathered significant attention recently [1, 2, 3, 4, 5]. The goal of these video captioning models is to produce a natural sentence that encapsulates the primary event in a short video. These models utilize datasets tailored to the described problem (e.g., MSR-VTT [6], VATEX [7]), where short video segments and their corresponding annotations in the form of captions are provided. Nevertheless, because real-world videos are often lengthy, untrimmed, and feature multiple simultaneous events alongside background content, the aforementioned single-sentence video captioning models typically produce lackluster and less informative descriptions. To tackle this more complex scenario, captioning approaches must both locate and describe the events occurring in long videos; this problem is known as dense video captioning [8, 9, 10, 11, 12].
@@
The majority of real-life videos encompass numerous events that may unfold simultaneously. For instance, in a video featuring “a waiter carrying food to a table” one may also observe another “individual eating and drinking” or a “woman sitting down”. Dense video captioning models are tasked with generating such descriptions based on captions for each of the events unfolding in the videos, precisely indicating the temporal instants of the start and end of these actions. The utility of such models is substantial, as they enable a spectrum of applications ranging from automatic caption generation systems to enhance search capabilities in multimedia databases, to solutions aiding individuals with visual impairments by providing a description of the unfolding events in the video.
@@
Figure 1: Above: Traditional models of dense video captioning work offline, accessing the whole video to generate the captions. Down: The live video captioning models must generate the captions for the video stream, in an online manner, ideally in real time, and working with partial observations of the video.
@@
The current state-of-the-art in dense video captioning primarily focuses on providing offline solutions. That is, as depicted in Figure 1, top part, all these models assume access to the entirety of the video for which they are tasked with generating captions. However, in this work we propose a paradigm shift which we term as Live Video Captioning (LVC), as illustrated in Figure 1, bottom part. In LVC, the challenge lies in generating dense captions in an online manner. This imposes significant new challenges on traditional offline models. Firstly, in an LVC scenario, it is not feasible to access the entire video to generate dense captions. LVC models must operate with partial observations of the video, ideally with video streams, continuously providing captions online as events unfold. This restriction is particularly detrimental to models requiring an action proposal generation phase, e.g. [13, 11, 14]. Secondly, LVC models must be able to anticipate actions. Operating with partial observations reduces the semantic information available compared to traditional offline models, making it more challenging to identify events. Furthermore, LVC solutions must implement temporal attention and filtering mechanisms to refine the caption predictions they generate. Finally, there is the challenge of real-time captioning inherent in live models. From an applicability standpoint, LVC solutions must function with both precision and speed, with both requirements sharing equal importance.
Interestingly, to the best of our knowledge, the problem of LVC has not been deeply investigated. In this work, we propose a specific LVC model, and our main contributions are as follows:
  1. 1.
    We provide a formal description of the live video captioning problem (Section III-A).
  2. 2.
    Alongside the problem description, we propose novel evaluation metrics specifically designed for the online scenario, demonstrating that offline metrics are not adequate (Section III-C).
  3. 3.
    We propose a specific model for LVC that integrates deformable transformers and a temporal filtering mechanism to generate dense captions in real-time over video streams (Section III-B).
  4. 4.
    Finally, in Section IV, we offer a comprehensive experimental evaluation on the ActivityNet Captions dataset, showcasing both the performance of our new model and the effectiveness of the new metrics. A comparison with state-of-the-art offline methods is also provided, highlighting their inadequacy for the proposed new online scenario.
II Related work
@@
We propose in this work an approach that addresses the problem of dense video captioning in an online fashion. This new problem is named as live video captioning (LVC). Ideally, in the LVC problem, the captions must be generated as soon as possible, by processing the video stream. This means that the models need to be adapted to work with partial observations of the video content, and, under this condition, produce dense captions as accurate as possible. Note that similar online approaches have been explored, for example, in the problem of action detection (e. g. [20, 21, 22, 23, 24, 25, 26]).
@@
For the dense video captioning problem, only, to the best of our knowledge, Hori et al. [27] have proposed a multi-modal captioning approach that uses a timing detector so that the captions can be generated in the early stages of an event-triggered video clip. This problem can be termed as early video captioning, where the target consists in evaluating the latency ratio needed to reach the same performance of an offline video captioning model for an event-triggered video. Similar simplified experimental setups where explored in the context of early event detection in video, e.g. [28, 29, 30]. Note that these problems are different from our live video captioning. We claim these simplified setups are not representative for practical applications, where occurrences of possibly many different actions need to be detected and a correct caption generated in an online manner, in long video recordings with widely varying content. When it comes to a live video captioning system, the model should be continuously processing the video stream, and, when necessary, producing dense video captions ideally in realtime. This necessitates precisely recognizing the current action at any point of its development. Furthermore, in order to complete the LVC task, one must distinguish the action from a range of negative input, such as background frames in which no pertinent actions are occurring.
@@
Overall, we propose a live video captioning model that parallelizes localization, selection, and captioning tasks within a single end-to-end model, based on deformable transformers, simplifying the process while ensuring the online generation of accurate and coherent captions. Our localization and captioning modules process the streaming video, producing dense captions that are enhanced with a filtering process.
III Live Video Captioning
III-A Live Video Captioning: problem formulation
@@
We define the problem of Live Video Captioning (LVC) as the process of obtaining dense captions for a video stream as soon as the video frames are available. Unlike the traditional video captioning model, which we refer to as offline video captioning, we do not have access to the entire video for analysis. Instead, LVC models can only access the content coming from a video stream up to the time instant $t$ to generate caption predictions for that instant. In other words, LVC models process the video in an online fashion, implying that the dense caption generation system has access to the current information of the video and past information, but never future information. Therefore, these systems are inherently causal.
We must assume that for LVC solutions, there is always information yet to be revealed. Caption predictions are made based on partial content of the video, in case it is already recorded. However, it is in the context of real-time video streams where LVC models gain special relevance. We can think of the following applications. For example, a surveillance camera, where we can never be certain about what will happen next, but we want to generate dense captions immediately after the information becomes accessible to our systems. We can also use the example of a robot equipped with a camera that generates descriptions of the scenes it is continuously seeing. It is not desirable to wait until the action is finished to have a correct caption prediction for it.
Therefore, the established properties for the LVC task in realistic scenarios are summarized as follows:
@@
    Timeliness: Captions must be generated as soon as the actions unfold, ideally in real-time.
@@
III-B Our Live Video Captioning Model
@@
As it is shown in Figure 2, we develop a deformable transformer model applied to the novel online dense video captioning problem. The deformable transformer model was introduced in [32] as an architecture to improve the performance of object detectors by attending to sparse spatial locations and incorporating multi-scale feature representations. Technically, our LVC integrates the deformable transformer-based architecture to make online caption predictions, taking temporal video segments of length $\Delta t$ as input. Given an input video stream $V_{i}=\left\{I_{1},I_{2},I_{3},\ldots\right\}$, we first split it in video segments $S_{i}$ of duration $\Delta t$ frames, hence $V_{i}=\left\{S_{1},S_{2},S_{3},\ldots\right\}$. Note that our LVC does not have to access the entire video, as required for offline systems such as [11, 33, 9, 10].
@@
These video snippets $S_{i}$ are the inputs for a deformable transformer model, with the corresponding encoder and decoder. The following operational scheme is followed from the introduction of the video segment to the model until the captions are obtained. First, our model extracts the features for each of the frames in the video segment $S_{i}$. Our LVC model can be integrated with any feature extractor. For our experiments, we used the TSP feature extractor [34]. To effectively use multi-scale features for forecasting events at different scales, we incorporate $L$ temporal convolutional layers with a stride of 2 and a kernel size of 3. This approach generates feature sequences at various resolutions, ranging from $T$ to $\frac{T}{2^{L}}$. These multi-scale features $\{\mathbf{x}^{l}\}_{l=1}^{L}$, combined with their positional embeddings, are input into the deformable transformer encoder, which captures frame-to-frame relationships across multiple scales.
The deformable transformer, as described in [32], is an encoder-decoder framework that utilizes multi-scale deformable attention (MSDeformAttn). For our set of multi-scale feature maps $\{\mathbf{x}^{l}\}_{l=1}^{L}$ where $\mathbf{x}^{l}\in\mathbb{R}^{C\times H_{l}\times W_{l}}$, a query element $\mathbf{q}_{j}$, and a normalized reference point $\mathbf{p}_{q}\in[0,1]^{2}$, MSDeformAttn produces a context vector via a weighted sum of $K\times L$ sampling points across the feature maps at $L$ scales:
@@
$\tilde{\mathbf{p}}_{jlk}=\phi_{l}(\mathbf{p}_{j})+\Delta\mathbf{p}_{jkl},$
In this equation, $\tilde{\mathbf{p}}_{jkl}$ and $A_{jkl}$ denote the position and attention weight of the $k$-th sampled key at the $l$-th scale for the $j$-th query element, respectively. $\mathbf{W}$ represents the projection matrix for key elements, and $\phi_{l}$ maps the normalized reference points into the feature map at the $l$-th level. The sampling offsets $\Delta\mathbf{p}_{jkl}$ are relative to $\phi_{l}(\mathbf{p}_{j})$. Both $A_{jkl}$ and $\Delta\mathbf{p}_{jkl}$ are determined through linear projection onto the query element. Overall, in the Deformable Transformer, self-attention modules in the Transformer encoder and cross-attention modules in the Transformer decoder are replaced with deformable attention modules.
@@
Localization head produces box prediction and binary classification for each event query. The box prediction task aims to determine the 2D relative offsets (center and length) of the ground-truth segment with respect to the reference point. Binary classification generates the foreground confidence for each event query. Both the box prediction and binary classification are carried out using multi-layer perceptrons. This process results in a set of tuples $[t_{j1},t_{jf},\alpha^{\rm loc}_{j}]_{j=1}^{N}$ that represent the detected events, with $t_{j1}$ and $t_{jf}$ the initial and final times, and where $\alpha^{\rm loc}_{j}$ is the localization confidence of the event query ${\tilde{{\mathbf{q}}}_{j}}$.
Instead of using a two-stage scheme, our LVC employs enhanced event query representations in parallel localization and captioning heads, allowing these two subtasks to be closely related. LVC directly produces a set of events with an appropriate size without relying on heuristic techniques to eliminate redundancy. Within our deformable transformer (see Figure 2), Event Queries are produced and introduced into the decoder. Each of these queries will result in a prediction for a caption. For all experiments, we use a total of 10 queries.
Our captioning head simply feeds ${\tilde{{\mathbf{q}}}_{j}}$ into a vanilla LSTM at each timestamp. The word $w_{jt}$ is predicted by a fully connected layer followed by a softmax activation over the hidden state $\mathbf{h}_{jt}$ of the LSTM.
@@
Confidence for each event query is calculated considering both location confidence and a modulated caption confidence to account for sentence length variability. During training, LVC generates a set of $N$ events with locations and captions. The Hungarian algorithm is used to match predicted events with ground truths, optimizing a cost function combining generalized IOU and focal loss. The total loss includes weighted sums of generalized IOU loss, classification loss, counting loss, and caption loss, computed across all transformer decoder layers.
While offline models for dense video captioning can access all available information by analyzing the entire video before generating the captions, our LVC model only has access to the information available in the video segment of duration $\Delta t$. This limitation can hinder the quality of the subtitles because a short video segment might not provide enough context.
To improve the quality of the captions, we have included a caption consolidation stage, as shown in Figure 3. The integrated transformer uses a query mechanism comprising 10 queries per video segment. Subsequently, the consolidation module accepts the predictions generated by these queries as input and orchestrates a voting mechanism, whereby the ultimate caption is determined by the highest frequency of occurrence among the generated predictions. In the voting process, we also take into account the score associated with the output head of the transformer, so that predictions with higher scores are favored.
@@
III-C Novel Evaluation Metric for Live Video Captioning: the Live Score

### v2

@@ Abstract
Dense video captioning involves detecting and describing events within video sequences. Traditional methods operate in an offline setting, assuming the entire video is available for analysis. In contrast, in this work we introduce a groundbreaking paradigm: Live Video Captioning (LVC), where captions must be generated for video streams in an online manner. This shift brings unique challenges, including processing partial observations of the events and the need for a temporal anticipation of the actions.
We formally define the novel problem of LVC and propose innovative evaluation metrics specifically designed for this online scenario, demonstrating their advantages over traditional metrics. To address the novel complexities of LVC, we present a new model that combines deformable transformers with temporal filtering, enabling effective captioning over video streams.
Extensive experiments on the ActivityNet Captions dataset validate the proposed approach, showcasing its superior performance in the LVC setting compared to state-of-the-art offline methods. To foster further research, we provide the results of our model and an evaluation toolkit with the new metrics integrated at: https://github.com/gramuah/lvc.
keywords:
@@
dense video captioning; online video analysis; transformers; deep learning; computer vision
1 Introduction
@@
As a growing field within video understanding, video captioning has gathered significant attention recently [1, 2, 3, 4, 5]. The goal of these video captioning models is to produce a natural sentence that encapsulates the primary event in a short video. These models utilize datasets tailored to the described problem (e.g., MSR-VTT [6], VATEX [7]), where short video segments and their corresponding annotations in the form of captions are provided. Nevertheless, because real-world videos are often lengthy, untrimmed, and feature multiple simultaneous events alongside background content, the aforementioned single-sentence video captioning models typically produce lackluster and less informative descriptions. To tackle this more complex scenario, captioning approaches must both locate and describe the events occurring in long videos; this problem is known as dense video captioning [1, 8, 9, 10, 11, 12].
@@
The majority of real-life videos encompass numerous events that may unfold simultaneously. For instance, in a video featuring “a waiter carrying food to a table” one may also observe another “individual eating and drinking” or a “woman sitting down”. Dense video captioning models must generate descriptions for each event unfolding in a video, precisely indicating the start and end times of each event.
@@
Figure 1: Above: Traditional models of dense video captioning work offline, accessing the whole video to generate the captions. Down: The live video captioning models must generate the captions for the video stream, in an online manner, and working with partial observations of the video.
@@
However, the current state-of-the-art in dense video captioning primarily offers offline solutions. As illustrated in the top part of Figure 1, these models assume access to the entire video before generating captions. In contrast, our work introduces a paradigm shift termed Live Video Captioning (LVC), depicted in the bottom part of Figure 1. In LVC, the challenge is to generate dense captions operating in an online manner. This introduces significant new challenges for traditional offline models. Firstly, in an LVC scenario, it is not feasible to access the entire video to generate dense captions. LVC models must work with partial video observations, ideally through video streams, continuously generating dense captions as events unfold. This restriction is particularly detrimental to models requiring an action proposal generation phase, e.g. [13, 11, 14]. Secondly, LVC models must be capable of anticipating actions. Working with partial observations limits the semantic information available compared to traditional offline models, making event identification more challenging. Lastly, LVC solutions need to incorporate temporal attention and filtering mechanisms to refine their caption predictions as the video streams evolve.
From an applicability standpoint, the LVC problem introduces new scenarios where live captioning is crucial, and traditional offline models would fail to perform effectively. LVC solutions must strike a balance between precision and speed, as both are equally vital. The utility of LVC models is substantial, enabling a wide range of applications, from content summarization and accessibility support in live broadcasts to online perception and decision-making in robotics and video surveillance. While the former paves the way for novel, accessible, and interactive multimedia experiences, the latter will be crucial for the next generation of autonomous navigation and human-robot interaction systems. These are, without question, scenarios where traditional offline video captioning models fall short.
Interestingly, to the best of our knowledge, the problem of Live Video Captioning (LVC) has not been extensively investigated. In this work, we make several key contributions. First, we offer a formal description of the LVC problem, detailed in Section 3.1. Second, we introduce a novel model for LVC that combines deformable transformers with a temporal filtering mechanism to generate dense captions over video streams, as described in Section 3.2. Third, we propose new evaluation metrics specifically designed for the online scenario, demonstrating that traditional offline metrics are inadequate, as discussed in Section 3.3. Finally, in Section 4, we provide a comprehensive experimental evaluation using the ActivityNet Captions dataset, showcasing both the performance of our new LVC model and the effectiveness of the new metrics. We also include a comparison with state-of-the-art offline methods, highlighting their limitations in addressing the new online scenario
2 Related work
@@
We propose in this work an approach that addresses the problem of dense video captioning in an online fashion. This new problem is named as live video captioning (LVC). Ideally, in the LVC problem, the captions must be generated as soon as possible, by processing the video stream. This means that the models need to be adapted to work with partial observations of the video content, and, under this condition, produce dense captions as accurate as possible. Note that other online approaches have been explored, for example, in the different problem of action detection (e. g. [20, 21, 22, 23, 24, 25, 26]).
@@
For the dense video captioning problem, only, to the best of our knowledge, Hori et al. [27] have proposed a multi-modal captioning approach that uses a timing detector so that the captions can be generated in the early stages of an event-triggered video clip. This problem can be termed as early video captioning, where the target consists in evaluating the latency ratio needed to reach the same performance of an offline video captioning model for an event-triggered video. Similar simplified experimental setups where explored in the context of early event detection in video, e.g. [28, 29, 30]. Note that these problems are different from our live video captioning. We claim these simplified setups are not representative for practical applications, where occurrences of possibly many different actions need to be detected and a correct caption generated in an online manner, in long video recordings with widely varying content. When it comes to a live video captioning system, the model should be continuously processing the video stream, and, when necessary, producing dense video captions. This necessitates precisely recognizing the current action at any point of its development. Furthermore, to complete the LVC task, one must distinguish the action from a range of negative inputs, such as background frames in which no pertinent actions are occurring.
@@
Overall, we propose a live video captioning model that parallelizes localization, selection, and captioning tasks within a single end-to-end model, based on deformable transformers [31, 32], simplifying the process while ensuring the online generation of accurate and coherent captions. Specifically, our approach is based on the deformable transformer of Zhu et al. [32], which is a model that was introduced as an architecture to improve the performance of object detectors by attending to sparse spatial locations and incorporating multi-scale feature representations. We have adapted this deformable transformer architecture to process video sequences for the novel problem of LVC. Our localization and captioning modules process the streaming video, producing online dense captions that are enhanced with an extra filtering process.
Similar transformers-based architectures were also used in offline dense video captioning models, e.g. [9, 10]. For instance, while both our model and [10] leverage deformable transformers, their architectural designs reflect fundamentally different objectives due to the contrasting demands of their respective tasks. In general, the main differences with respect the previous offline dense video captioning models [9, 10], that also use transformer-based architectures are as follows. First, they assume full access to the entire video sequence during both training and inference. This enables the models to perform holistic event localization and captioning by leveraging global temporal information. In contrast, our LVC model processes partial, sequentially available video segments under causal constraints, where no future frames are accessible. To accommodate this, our deformable transformer is modified to focus solely on the features within the current segment and refine predictions dynamically as new frames arrive. Second, in the offline dense video captioning setting, a multi-scale deformable attention can be used to process densely extracted frame features from the entire video. This attention mechanism ensures that the model captures relationships between temporally distant events. In our LVC model, we adapt the multi-scale deformable attention mechanism to operate on temporally truncated input streams. For instance, our event queries are tailored for online processing, designed to anticipate and refine event boundaries dynamically as frames are processed. This iterative refinement aligns with the causal nature of the LVC task, allowing the model to balance precision with latency. We also add a temporal filtering mechanism, integrated with the learnable event queries, to dynamically generate the dense captions for the events within the limited temporal context available, enabling precise and actionable online predictions. These distinctions highlight how our use of deformable transformers is uniquely adapted to the challenges of LVC.
In table 1 we provide a global comparison of the main models for dense video captioning. Note that only our approach has been specifically designed for the novel LVC problem.
Work Joint Loc. & Cap. LVC Contributions
[13] First work; Proposals module + attention-based LSTM for caption generation.
[15, 16] Enriched event representations by context modeling.
[12] Incorporate event-level relationship.
[17, 9] Use of multi-modal feature fusion.
[18] Use of proxy task based on language rewards to enhance the localization.
[10] Inter-task interactions with deformable transformers.
[19] Differential masking for linking the gradient flow from captioning loss to proposals’ boundaries.
LVC (Ours) Deformable transformers + temporal filtering for video streams.
Table 1: Review of other models for the problem of dense video captioning. We organize the literature based on whether the models address the problems of localization and captioning jointly, and whether they can handle the proposed LVC scenario or operate offline. We also summarize some of the main contributions of the works.
3 Live Video Captioning
3.1 Live Video Captioning: problem formulation
@@
We define the problem of Live Video Captioning (LVC) as the process of obtaining dense captions for a video stream as soon as the video frames are available. Unlike the traditional video captioning scenario, which we refer to as offline video captioning, we do not have access to the entire video for analysis. Instead, LVC models can only access the content coming from a video stream up to the time instant $t$ to generate caption predictions for that instant. In other words, LVC models process the video in an online fashion, implying that the dense caption generation system has access to the current information of the video and past information, but never future information. Therefore, these systems are inherently causal.
LVC models face unique challenges when compared to traditional offline captioning systems. In offline systems, the model has access to the entire video, allowing for accurate temporal segmentation and caption generation after full observation. In contrast, LVC models must generate captions in an online fashion, working with partial, often incomplete frames, and make predictions with limited information. This online constraint introduces several challenges: (i) accurate temporal segmentation with minimal context, (ii) the need for anticipation of actions before they are fully visible, and (iii) ensuring that caption generation remains both fast and precise.
We must assume that for LVC solutions, there is always information yet to be revealed. Caption predictions are made based on partial content of the video, in case it is already recorded. However, it is in the context of live video streams where LVC models gain special relevance. We can think of the following applications and problems. For instance, consider a surveillance camera monitoring a busy public space, where the flow of events is unpredictable. In such a scenario, we can never be certain about what might occur next—a sudden altercation, a person leaving an unattended bag, or an individual entering a restricted area. However, it is crucial to generate dense captions as soon as the visual information is available to the system. This capability enables timely detection and response to critical events, enhancing safety and situational awareness. Another illustrative example is a robot equipped with a camera, tasked with generating detailed descriptions of the scenes it encounters while navigating a dynamic environment, such as a bustling factory floor or a crowded shopping mall. The robot might describe activities like workers assembling parts, customers interacting with products, or obstacles in its path. This continuous scene interpretation is vital for tasks like human-robot interaction, navigation, and situational understanding. Any traditional offline model for dense video captioning would not be able to generate these descriptions, as these models have been trained to operate only when the entire video is available. For instance, models that rely on action proposals (e. g. [13, 15, 16]) would be severely affected. In the novel LVC paradigm, it is not desirable to wait until the action is finished to have a correct caption prediction for it.
Therefore, a critical advantage of Live Video Captioning (LVC) is its ability to anticipate actions and provide temporally precise localization of events, a limitation commonly observed in offline dense video captioning methods. For instance, in a video where a person begins to raise a glass to drink, followed by a second person setting a plate on the table, offline methods would generate overlapping captions like ’[2.0s-6.0s] A person moves’ and ’[5.0s-9.0s] A person interacts with an object,’ failing to distinguish between the two actions clearly or provide precise temporal boundaries. By contrast, LVC models would produce ’[2.0s-4.0s] A person raises a glass’ and ’[4.0s-6.0s] Another person places a plate on the table,’ with non-overlapping, accurate temporal segmentation. This novel scenario not only allows for early recognition of ongoing events but also improves temporal granularity, enabling a clearer distinction between simultaneous or sequential actions. These capabilities are vital for applications requiring immediate and accurate action detection, such as the monitoring systems or interactive robotics described.
In the new context of LVC, we proceed to define the fundamental properties that characterize it, which are as follows:
@@
    Timeliness: Captions must be generated as soon as the actions unfold.
@@
3.2 Our Live Video Captioning Model
@@
As it is shown in Figure 2, we develop a deformable transformer model applied to the novel online dense video captioning problem. The deformable transformer model was introduced in [32] as an architecture to improve the performance of object detectors by attending to sparse spatial locations and incorporating multi-scale feature representations. Technically, our LVC model integrates the deformable transformer-based architecture to construct online caption predictions, taking temporal video segments of length $\Delta t$ as inputs. Given an input video stream $V_{i}=\left\{I_{1},I_{2},I_{3},\ldots\right\}$, we first split it in video segments $S_{i}$ of duration $\Delta t$, hence $V_{i}=\left\{S_{1},S_{2},S_{3},\ldots\right\}$. Note that our LVC does not have to access the entire video, as required for offline systems such as [11, 33, 9, 10].
@@
These video snippets $S_{i}$ are the inputs for a deformable transformer model, with the corresponding encoder and decoder. The following operational scheme is followed from the introduction of the video segment to the model until the captions are obtained. First, our model extracts the features for each of the frames in the video segment $S_{i}$.
For our goal of generating dense captions for the actions and events occurring in videos, we have chosen to use a pre-trained model for action recognition. In particular, for our experiments, we used the Temporally-Sensitive Pretraining (TSP) feature extractor [34]. However, our LVC model can be integrated with any action recognition feature extractor. We then rescale the temporal dimension of the feature maps obtained to a fixed length $T$ using interpolation. To effectively use multi-scale features for forecasting events at different scales, we incorporate $L$ temporal convolutional layers with a stride of 2 and a kernel size of 3. This approach generates feature sequences at various resolutions, ranging from $T$ to $\frac{T}{2^{L}}$. These multi-scale features $\{\mathbf{x}^{l}\}_{l=1}^{L}$, combined with their positional embeddings, are input into the deformable transformer encoder, which captures frame-to-frame relationships across multiple scales.
Note that, accurate captioning for videos with multiple simultaneous events relies heavily on temporal information to segment overlapping actions and on contextual information to maintain coherence. Our model incorporates these elements through the proposed deformable transformer-based architecture. Temporal information is leveraged via multi-scale temporal convolutional layers that enable the extraction of features at varying time resolutions, capturing the precise start and end of events. Contextual relationships are modeled using the deformable attention mechanism, which dynamically attends to relevant temporal regions, allowing the system to refine predictions based on localized action patterns. These components work in tandem to disambiguate concurrent events and generate distinct, temporally aligned captions, even in complex multi-event scenarios. Moreover, our model integrates the multi-scale temporal convolutional layers to aggregate features across consecutive frames, enabling the detection of temporal patterns that extend beyond individual frame boundaries.
Our deformable transformer, as described in [32], is an encoder-decoder framework that utilizes multi-scale deformable attention (MSDeformAttn). For our set of multi-scale feature maps $\{\mathbf{x}^{l}\}_{l=1}^{L}$ where $\mathbf{x}^{l}\in\mathbb{R}^{C\times H_{l}\times W_{l}}$, a query element $\mathbf{q}_{j}$, and a normalized reference point $\mathbf{p}_{q}\in[0,1]^{2}$, MSDeformAttn produces a context vector via a weighted sum of $K\times L$ sampling points across the feature maps at $L$ scales:
@@
$\tilde{\mathbf{p}}_{jlk}=\phi_{l}(\mathbf{p}_{j})+\Delta\mathbf{p}_{jkl}.$
In this equation, $\tilde{\mathbf{p}}_{jkl}$ and $A_{jkl}$ denote the position and attention weight of the $k$-th sampled key at the $l$-th scale for the $j$-th query element, respectively. $\mathbf{W}$ represents the projection matrix for key elements, and $\phi_{l}$ maps the normalized reference points into the feature map at the $l$-th level. The sampling offsets $\Delta\mathbf{p}_{jkl}$ are relative to $\phi_{l}(\mathbf{p}_{j})$. Both $A_{jkl}$ and $\Delta\mathbf{p}_{jkl}$ are determined through linear projection onto the query element. Overall, in our deformable transformer, self-attention modules in the transformer encoder and cross-attention modules in the transformer decoder are replaced with deformable attention modules.
@@
Our localization head produces a box prediction and a binary classification for each event query. The box prediction task aims to determine the 2D relative offsets (center and length) of the ground-truth segment with respect to the reference point. Binary classification generates the foreground confidence for each event query. Both the box prediction and binary classification are carried out using multi-layer perceptrons. This process results in a set of tuples $[t_{j1},t_{jf},\alpha^{\rm loc}_{j}]_{j=1}^{N}$ that represent the detected events, with $t_{j1}$ and $t_{jf}$ the initial and final times, and where $\alpha^{\rm loc}_{j}$ is the localization confidence of the event query ${\tilde{{\mathbf{q}}}_{j}}$.
Instead of using a two-stage scheme, our LVC employs enhanced event query representations in parallel localization and captioning heads, allowing these two subtasks to be closely related. LVC directly produces a set of events with an appropriate size without relying on heuristic techniques to eliminate redundancy. Within our deformable transformer (see Figure 2), Event Queries are produced and introduced into the decoder. Each of these queries will result in a prediction for a caption. For all experiments, we use a total of 10 queries, although in the ablation studies we evaluate the impact of this parameter.
Our captioning head feeds ${\tilde{{\mathbf{q}}}_{j}}$ into a vanilla LSTM at each timestamp. The word $w_{jt}$ is predicted by a fully connected layer followed by a softmax activation over the hidden state $\mathbf{h}_{jt}$ of the LSTM. Note that this LSTM operates only within the video segment of duration $\Delta t$ that our LVC model is analyzing.
@@
Confidence for each event query is calculated by combining two components: location confidence and a modulated caption confidence. The location confidence evaluates the spatial accuracy of the event, while the modulated caption confidence incorporates adjustments for variability in sentence length. During training, our LVC model generates a set of $N$ predicted events, each comprising a spatial location and an associated caption. To associate these predictions with the ground truth, the Hungarian algorithm is employed, solving a bipartite matching problem that optimizes a comprehensive cost function. This cost function integrates two primary components: the Generalized Intersection over Union (GIoU) loss [35] for spatial alignment and a focal loss [36] for classification accuracy.
Our total training loss, $\mathcal{L}_{\text{total}}$, is computed as a weighted combination of several loss terms: $\mathcal{L}_{\text{total}}=\lambda_{\text{GIoU}}\mathcal{L}_{\text{GIoU}}+\lambda_{\text{cls}}\mathcal{L}_{\text{cls}}+\lambda_{\text{count}}\mathcal{L}_{\text{count}}+\lambda_{\text{cap}}\mathcal{L}_{\text{cap}}$, where $\mathcal{L}_{\text{GIoU}}$, $\mathcal{L}_{\text{cls}}$, $\mathcal{L}_{\text{count}}$, and $\mathcal{L}_{\text{cap}}$ represent the GIoU loss, classification loss (i.e. focal loss), counting loss, and caption loss, respectively. Our counting loss is the cross-entropy loss between the estimated count distribution and the one in the ground-truth. For the captioning loss we use the cross-entropy computed between the word probability estimated by our model and the ground-truth normalized by the length of the caption. The hyperparameters $\lambda_{\text{GIoU}}$, $\lambda_{\text{cls}}$, $\lambda_{\text{count}}$, and $\lambda_{\text{cap}}$ control the relative contributions of each loss term. For our experiments we use $\lambda_{\text{GIoU}}=2$, and $\lambda_{\text{cls}}=\lambda_{\text{count}}=\lambda_{\text{cap}}=1$.
While offline models for dense video captioning can access all available information by analyzing the entire video before generating the captions, our LVC model only has access to the information available in the video segment of duration $\Delta t$. This limitation can hinder the quality of the subtitles because a short video segment might not provide enough context. In our LVC model, captions are generated in an online manner based exclusively on the video data available up to the current time step. Once a caption is produced for a given video segment, it remains immutable and is not revised as new frames are processed. This design choice enforces the causality constraints essential to the LVC task, ensuring that each caption accurately reflects the observed content at the moment of generation without relying on future information. This approach distinctly differentiates LVC from traditional offline methods, which have access to the entire video sequence and can even refine captions retrospectively.
Therefore, to improve the quality of the captions, we have included a temporal filtering for the caption consolidation, as shown in Figure 3. First, our deformable transformer uses a query mechanism comprising $N$ queries per video segment of duration $\Delta t$. Subsequently, our filtering module accepts the predictions generated by these event queries, $\{\mathbf{q}_{j}\}_{j=1}^{N}$, as input. Each query produces a predicted caption $c_{j}$, an associated temporal segment $[t_{j1},t_{jf}]$, and a confidence score $\alpha_{j}$, resulting in a set of tuples $\{(t_{j1},t_{jf},c_{j},\alpha_{j})\}_{j=1}^{N}$. Considering the temporal overlap between the predictions, the consolidation module orchestrates a voting mechanism to determine the final captions for the given video segment. Specifically, the temporal filtering we propose consists in a module that evaluates the frequency of occurrence of each overlapped predicted caption among the set of predictions. Let $f(c)$ represent this frequency for a particular caption $c$, and $w(c)$ its maximum score. The ultimate caption $c_{*}$ for the set of temporally overlapped predictions is selected as the one with the highest frequency: $c_{*}=\arg\max_{c\in\{c_{1},\dots,c_{N}\}}f(c)$. Being the final tuple for the prediction as follows: $\{(t_{*1},t_{*f},c_{*},w(c_{*}))\}$.
@@
3.3 Novel Evaluation Metric for Live Video Captioning: the Live Score